The equal sides of an isosceles triangle are 8 cm long and the radius of its cir
cum circle is 5 cm calculate the length of its third side

Answer by ewatrrr(24785)


x^2 + y^2 = 25
x^2 + (5+y)^2 = 64
(25-y^2 ) + 25 + 10y + y^2 = 64
10y = 14
y = 1.4
x^2 + (1.4)^2 = 25
x = 4.8
length of 3rd side is 9.6 cm